# The physical basis of teleological evolutions

### v11

@@ Abstract
We show that the sheer existence of a quantum computational speedup logically implies the mutually exclusive $or$ of well-defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. Involving only the measurements of commuting observables, quantum algorithms are submitted to the same (classical) logic and should therefore host the quantum superposition of the causal loops in question. However, their ordinary quantum description does not and , being causal in character, cannot describe causal loops. It must therefore be incomplete and is in fact completed by time-symmetrizing it. This leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered but replaces its single causal process by the quantum superposition of the causal loops in question. In their completed quantum description, quantum algorithms respond to the pre-scientific notion of teleological evolution, that is, an evolution toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again). Once applied to the evolutions of the living, this notion was dismissed with the advent of modern science for the alleged absence of a physical basis. We show that, under a (hypothetical) quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
@@ 1 Rationale
This work proposes a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the very goal it will produce in the future. It is, of course, a puzzling notion since it implies a causal loop. For more than two millennia, it has been applied by natural philosophy first to the functional structure of the living (Aristotle: teeth were created for the purpose of chewing) and then, with the advent of the notion of evolution of the living, to it. It was then dismissed by modern science because of its alleged lack of physical basis.
What we believe should change this (centuries-old) position of science is the advent of quantum mechanics and, more specifically, the relatively recent discovery of the quantum computational speedup. In fact, we will show that quantum algorithms are scientific examples of teleological evolutions: it is their teleological character that yields their quantum speedup over their classical counterparts.
This is not just the rehabilitation by science of an abandoned philosophical notion. Once represented in a scientific way, the latter fires back at science by challenging its current (long standing) way of relying on causality, that is, on unitary evolutions in which causality always goes in the same time direction. This is not only a new interpretation of the foundations of quantum mechanics, in fact we will see that it provides a variety of unifying quantitative explanations of phenomena that would be otherwise unattainable.
Dealing with causal loops in the nonrelativistic quantum framework, the work is of course unconventional. To explain in sufficient detail from the beginning an entirely different approach to quantum computation, to the usual “Introduction” we have preferred a longer “Rationale”. The latter is divided into two parts. The first provides a physical basis to the teleological character of quantum algorithms and, more in general, the building up of quantum correlation (in quantum algorithms, between the setting and the solution of the problem). The second explores the possibility that the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living for which the teleological notion was originally conceived.
@@ 1.1 The physical basis of the teleological notion in the quantum domain
We must begin with an example of quantum speedup. The context, of course, is problem-solving. For example, Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem; if it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case.
Let us show, by the way very simply, that always solving the problem of locating the ball by opening only one drawer logically implies the mutually exclusive $or$ of well defined causal loops. In fact, it is logically equivalent to the problem-solver knowing in advance, before she begins her problem-solving action, that the ball is in one of two given drawers and benefiting of this knowledge to solve the problem by opening just one drawer. Indeed, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer. Note that this also means knowing in advance half of the information that specifies the solution of the problem she will produce and measure in the future.
@@
The above immediately generalizes to any number, $N$, of drawers. Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case222 The four drawer instance of Grover algorithm gives the solution with certainty; with more than four drawers, by “Grover algorithm” we will always mean Long’s version of it $\left[2\right]$, which always yields the solution with certainty. . Also now, this quantum speedup can be logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers. She can then solve the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
Note the simplicity of the above argument. It is so simple that it runs the risk of being overlooked. Instead, we argue, it should be of fundamental importance. Whether causal loops are physical or unphysical is a controversial subject. Of course, they are never used in mainstream quantum mechanics. According to the argument in question, instead, they must be implicit in Grover quantum search algorithm. Moreover, as we will see, they are necessarily implicit in any quantum algorithm that yields a quantum speedup over what is classically possible.
We should now move from logic to physics. Of course, in some way, the causal loops in question should be implicit in the quantum algorithm; it is then important to see how they are physically hosted in it. Since quantum algorithms only involve the measurements of commuting observables (the problem-setting and the corresponding solution), they are submitted to the same classical logic of the logical implication in question. Therefore, given that the mutually exclusive $or$ of possibilities in the quantum world becomes their quantum superposition, we should expect that a complete quantum description of quantum algorithms hosts the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by their quantum speedup.
@@
Now, of course, the current quantum description of quantum algorithms does not describe any causal loop. Furthermore, being the customary “causal” quantum description of mainstream quantum mechanics, where causality always goes in the same time-direction, it could not. Consistently with the present line of thought, we must conclude that it is incomplete.
@@
The problem is then how to complete it. Obviously, to start with, we should choose the kind of quantum description of quantum algorithms that has to be completed. This cannot be their usual quantum description, which is itself incomplete in a trivial way. It is in fact limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It thus lacks the initial measurement: of course, the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement.
@@
Naturally, this problem is simply amended. It suffices to extend the quantum description of the process of solving the problem to the upstream process of setting the problem. An initial measurement randomly selects a problem-setting out of a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, the unitary computation of the corresponding solution and the final measurement of the latter follow; for simplicity, we will always assume that the computation of the solution yields it with certainty. In view of what will follow, note that, as a consequence, the unitary transformation goes from the initial to the final measurement outcome despite comprising the final measurement. Thus, in particular, the process connecting the two measurement outcomes is time-reversible.
@@
We require that the determination of the problem-setting and thus that of the corresponding solution, instead of being entirely performed by the initial measurement of the problem-setting, evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. Since there is a plurality of mutually exclusive ways of evenly sharing, at the end we should take their quantum superposition.
Note that, since the unitary evolution between the two measurement outcomes does not change the problem-setting selected by the initial measurement, the latter measurement, or any part of it, can be postponed at will, thus also making it simultaneous with the final measurement. Of course, this allows to share the selection between initial and final measurements. However, the fact that the sharing is possible of course does not say why we should assume it. The answer is that this is the only way of generating the causal loops that the quantum algorithm must host. An obvious question is how the initial measurement of the problem-setting knows that there will be a future measurement of the corresponding solution with which to share the selection of the pair. Some form of quantum nonlocality should be involved. We will soon see that this is exactly the case.
In each time-symmetrization instance, the state selected by the initial measurement (a suitable partition of the quantum superposition of all the problem-settings) unitarily propagates forward in time until becoming the state immediately before the final measurement (a quantum superposition of the problem-settings selected by the initial measurement each tensor product the corresponding solution). Time-symmetrically, the state selected by the final measurement (a single problem-setting tensor product the corresponding solution) propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous outcome of the initial measurement.
Of course, propagating backward in time the outcome of the final measurement is completely unconventional. However, as we will soon see, it is mandatory to complete the quantum description of the quantum algorithm in a way that describes the causal loops that it must host.
@@
Let us call for short the above sequence of operations the time-symmetrization of the ordinary quantum description of the quantum algorithm, because it is indeed a time-symmetrization of the time-reversible quantum process between the initial and final one-to-one correlated measurement outcomes. The time-symmetrized quantum algorithm is the backward in time propagation, which inherits both selections. Note that, however, the complete quantum description of the quantum algorithm is the description of the entire time-symmetrization process, not only of its end result.
@@
(i) each time symmetrization instance consists of a time-forward propagation (by a unitary transformation) going from the outcome of the initial measurement to the outcome of the final measurement and a time-backward propagation (by the inverse of the previous unitary transformation) going from the outcome of the latter measurement until replacing the previous outcome of the initial measurement. One can see that this describes a causal loop that changes its own past.
(ii) In each time-symmetrization instance, the outcome of the final measurement – a single problem-setting tensor product the corresponding solution – coincides with that of the ordinary quantum description of the quantum algorithm. As a consequence, mathematically (i.e. disregarding the time-direction of causality), its backward in time propagation, by the inverse of the time-forward unitary transformation, is the unitary part of the ordinary quantum description of the quantum algorithm back again. Of course, this holds for each time-symmetrization instance and consequently for their quantum superposition.
We can see how this time-symmetrization completes the ordinary (causal) quantum description of the quantum algorithm. First, it leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered: it is indeed the end result of the time-symmetrization process, that is, its backward in time propagation. Second, it changes the description of the behavior of causality along that unitary part: the single causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
@@
Let us explain the reason for the specification “as viewed by the problem-solver”. The ordinary quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. It cannot be with respect to the problem-solver. In fact, it would tell her, before she begins her problem-solving action, the problem-setting selected by the initial measurement and thus the solution of the problem (here, they are both the number of the drawer with the ball). Therefore, we should relativize with respect to the problem-solver the causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, by hiding from her the problem-setting selected by the initial measurement. See Subsection 2.1.2. for further detail.
Also note that the time-symmetrization we are dealing with is the only possible way of completing the ordinary quantum description of the quantum algorithm in a way that describes the causal loops logically implied by the sheer existence of its quantum speedup. In fact, to describe a causal loop (any one) we must properly share the selection of the problem-setting between initial and final measurements with the part of the selection performed by the initial measurement propagating forward in time and that performed by the final measurement backward in time. To obtain exactly the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, this sharing must be even. This exactly identifies the time-symmetrization we have performed. It also shows that this time-symmetrization is mandatory: it is the only way of generating the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup and that the quantum algorithms that delivers it must host.
Further below we will show that all this holds for a large class of quantum algorithms, those that solve oracle problems. Note that, so completed, the quantum description of quantum algorithms says much more than their ordinary quantum description; it explains the reason for the quantum speedup in the following unified quantitative way:
@@
Also note that, in mainstream quantum computing science, which of course relies on the dominant notion of causal evolution, the only way of knowing the number of computation steps needed to solve a given oracle problem in an optimal quantum way is finding a quantum algorithm that does that, naturally an object of great ingenuity and uncertain achievement.
We call the above statement in italics the advanced knowledge rule. This rule had already been achieved in the retrocausal approach to the quantum speedup of ourselves and coauthors $\left[3-7\right]$. In $\left[6\right]$, we have shown that it quantitatively explains the quantum speedup of all the major quantum algorithms. However, in it, the time-symmetrization of the ordinary quantum description of the quantum algorithm was justified only by the fact that it turns out to produce, indeed, a unified quantitative description of the quantum speedups.
Now, this time-symmetrization has been justified in a different way. Indeed, we started from the fact – we think inescapable – that the sheer existence of a quantum speedup over what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops that must be implicit in the quantum algorithm that delivers it. We have seen that the only possible way of generating the quantum description of the causal loops in question is by time-symmetrizing the ordinary quantum description of the quantum algorithm.
Summarizing, this time we would have demonstrated that causal loops, of the present form, are commonplace in nonrelativistic quantum mechanics, at least as common as quantum algorithms are. As causal loops are not yet accepted by mainstream quantum mechanics, this should be a significant advancement.
In Section 2.2 we will see that all the above is not confined to quantum algorithms but generalizes in the following rule:
@@
In the first place, the rule in question says that the building up of quantum correlation between two initially unrelated observables can always occur with quantum speedup. This is the case of quantum algorithms, but we will see that it is not limited to them. By the way, the notion of advanced knowledge (on the part of the problem-solver) of half of the information about the solution could lose any meaning outside problem-solving. However, note that it can be replaced by that of a corresponding reduction of the dimension of the Hilbert space in which the search for one-to-one correlation between the two observables takes place.
In the second place, the causal loop rule of quantum correlation also applies to the case that the unitary evolution between the two one-to-one correlated measurement outcomes does not build up any quantum correlation because the two observables that will be measured are maximally entangled to start with. This is the case of quantum nonlocality, where there is one-to-one quantum correlation between the outcomes of measuring (in the same basis) two maximally entangled observables in a state of spatial separation between the two respective subsystems.
The quantum process that generates quantum nonlocality can be seen as a unitary evolution between two one-to-one correlated measurement outcomes that goes, first, backward in time from the outcome of the first measurement to the time when the two subsystems were not yet spatially separate and then (in causal order) forward in time to the outcome of the second measurement. Of course, the eigenvalue selected by either measurement is not changed by the unitary process of space separation between the two subsystems.
The causal loops generated by time-symmetrizing the ordinary quantum description of the process in question also locally go from the initial to the final measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell the not yet measured observable the state to assume in view of its future measurement $\left[7\right]$. Therefore, these causal loops exactly play the role of the hidden variables that Einstein and the others $\left[8\right]$ envisaged to locally explain quantum nonlocality. Obviously, this also says that the incompleteness that we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and the others ascribed to that of quantum nonlocality. Of course, in both cases we are dealing with the incompleteness of the quantum description of a quantum correlation.
@@
All the above also highlights the teleological character of quantum algorithms, indeed their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again). We can say that, in the quantum world, we have found the missing physical basis for the notion of teleological evolution.
@@
This, naturally, raises the question whether the physical basis in question also applies to the evolutions of the living for which the notion of teleological evolution was originally conceived. The second part of the work addresses this question.
@@
Incidentally, being thought without physical basis, and because of its finalistic character, the teleological view has been classified among the beliefs. This, however, should not conceal the fact that it was motivated by obvious empirical reasons. Indeed, it fits the functional structures and the evolutions of the living much better than the causal view. Of course, this must be why it was conceived in the first place and, in relatively recent times, whether to abandon it has long been debated among evolutionary biologists.
@@
Of course, the question of whether the physical basis of the teleological character of quantum algorithms also applies to natural evolutions immediately raises a seemingly insurmountable problem. Quantum algorithms belong to the microscopic world; their implementations deal with a rather limited number of quantum bits and durations of a few tens microseconds. As things are now, larger sizes or durations soon bring to quantum decoherence and the process becomes classical. The evolutions of the living, instead, can be as large as the Earth and have durations of billions of years. They are therefore among the furthest things from the quantum processes of the microscopic world to which quantum algorithms belong.
@@
Of course, we are not discussing whether the chlorophyll cycle or orientation in migratory birds is quantum, which could be a possibility. Here, the very enormity of the distance forces us to avoid any attempt to scale up toward the macroscopic the quantum features of the microscopic world.

### v12

@@ Abstract
We show that the sheer existence of a quantum computational speedup logically implies the mutually exclusive $or$ of well-defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. Involving only the measurements of commuting observables, quantum algorithms are submitted to the same (classical) logic and should therefore host the quantum superposition of the causal loops in question. However, their ordinary quantum description does not and, being causal in character, cannot describe causal loops. It must therefore be incomplete and is in fact completed by time-symmetrizing it. This leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered but changes the behavior of causality along it. The single causal process of the ordinary quantum superposition is replaced by the quantum superposition of the causal loops in question. In their completed quantum description, quantum algorithms respond to the pre-scientific notion of teleological evolution, that is, an evolution toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again). Once applied to the evolutions of the living, this notion was dismissed with the advent of modern science for the alleged absence of a physical basis. We show that, under a quantum cosmological interpretation of the Anthropic Principle, the same quantum superposition of causal loops underlying the teleological character of quantum algorithms becomes the missing physical basis of the teleological character of natural evolutions.
@@ 1 Rationale
This work proposes a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the very goal it will produce in the future. It is, of course, a puzzling notion since it implies a causal loop. For more than two millennia, it was applied by natural philosophy first to the functional structure of the living (Aristotle: teeth were created for the purpose of chewing) and then, with the advent of the notion of evolution of the living, to these evolutions. It was then dismissed by modern science because of its alleged lack of physical basis.
What we believe should change this (centuries-old) position of science is the advent of quantum mechanics and, more specifically, the relatively recent discovery of the quantum computational speedup. In fact, we will show that quantum algorithms are scientific examples of teleological evolutions: it is their teleological character that yields their quantum computational speedup over the corresponding classical algorithms.
This is not just the rehabilitation by science of an abandoned philosophical notion. Once represented in a scientific way, the teleological notion fires back at mainstream science by challenging its current rejection of the notion of retrocausality.
To explain in sufficient detail from the beginning the present unconventional approach to quantum computation, to the usual “Introduction” we have preferred a longer “Rationale”. The latter is divided into two parts. The first provides a physical basis to the teleological character of quantum algorithms and, more in general, the building up of quantum correlation (in quantum algorithms, between the setting and the solution of the problem). The second discusses the possibility that the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living.
@@ 1.1 The physical basis of the teleological notion in the quantum domain
We must begin with an example of quantum speedup. The context, of course, is problem-solving. We consider the following very simple problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem; if it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the optimal quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case.
Let us show, by the way very simply, that always solving the problem of locating the ball by opening only one drawer logically implies the mutually exclusive $or$ of well defined causal loops.
First, it is logically equivalent to the problem-solver always knowing in advance, before she begins her problem-solving action, that the ball is in one of two given drawers and benefiting of this knowledge to solve the problem by opening just one drawer. Indeed, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer. Note that this also means knowing in advance half of the information that specifies the solution of the problem she will produce and measure in the future.
@@
The above immediately generalizes to any number, $N$, of drawers. Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case222 The four drawer instance of Grover algorithm gives the solution with certainty; with more than four drawers, by “Grover algorithm” we will always mean Long’s version of it $\left[2\right]$, which always yields the solution with certainty. . Also now, this quantum speedup can be logically justified by the fact that Alice always knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers. She can then solve the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
Note the simplicity of the above argument. It is so simple that it runs the risk of being overlooked. Instead, we argue, it should be of fundamental importance. Whether causal loops are physical or unphysical is a controversial subject. Of course, they are never used in mainstream quantum mechanics. According to the argument in question, instead, they must be implicit in Grover quantum search algorithm. Moreover, as we shall see, they are necessarily implicit in any quantum algorithm that yields a quantum speedup over what is classically possible.
@@
Let us now move from logic to physics. Of course, the current quantum description of quantum algorithms does not describe any causal loop. Furthermore, being the customary “causal” quantum description of evolutions, where causality always goes in the same time-direction, it could not. Consistently with the present line of thought, we must conclude that this quantum description is incomplete.
This is, of course, a bold claim since, as it is, the quantum description has worked amazingly well for exactly a century to date. However, the boldness of the claim can be mitigated in two ways:
(i) First, it is a claim with an illustrious precedent, although apparently forgotten by the scientific community. We mean the 1935 argument of Einstein et al on the incompleteness of the quantum description $\left[3\right]$. We will see that the incompleteness of the quantum description of quantum algorithms is exactly the same lamented by Einstein et al for quantum nonlocality; moreover, the way to complete the quantum description of quantum algorithms also applies to quantum nonlocality and provides the hidden variables needed in both cases.
(ii) Second, let us distinguish between the mathematical and the physical description of a quantum process. The completion of the quantum description we will propose leaves the unitary part of the quantum algorithm mathematically unchanged but changes the description of the behavior of causality along it. The single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutual exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
@@
We show how this result is reached. Here we provide conceptually what will be formalized in Section 2.. Obviously, to start with, we should choose the kind of quantum description of quantum algorithms to complete. This is said because it cannot be their usual quantum description, which is itself trivially incomplete. Indeed, it is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It thus lacks the initial measurement. Of course, the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement.
@@
Naturally, this problem is simply amended. It suffices to extend the quantum description of the process of solving the problem to the upstream process of setting the problem. An initial measurement randomly selects a problem-setting (e. g. the number of the drawer with the ball) out of a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, the unitary computation of the corresponding solution and the final measurement of the latter follow; for simplicity, we will always assume that the computation of the solution yields it with certainty. In view of what will follow, we note that, as a consequence, the unitary transformation goes from the initial to the final measurement outcome despite comprising the final measurement; thus, in particular, the process connecting the two measurement outcomes is time-reversible.
@@
We require that the determination of the problem-setting and thus that of the corresponding solution, instead of being entirely ascribed to the initial measurement of the problem-setting, is evenly shared between the initial measurement of the problem-setting and the final measurement of the corresponding solution. Since there is a plurality of mutually exclusive ways of evenly sharing, at the end we should take their quantum superposition.
Note that, since the unitary evolution between the two measurement outcomes (i.e., the computation of the solution) never changes the problem-setting selected by the initial measurement, the latter measurement, or any part of it, can be postponed at will, thus also making it simultaneous with the final measurement. Of course, this allows the sharing of the selection in question but does not justify the fact of assuming it. As we shall see, the justification is that it is the only way to generate the causal loops the quantum algorithm must host.
In each time-symmetrization instance, the state selected by the initial measurement (not a single problem-setting but a suitable partition of the quantum superposition of all the problem-settings) unitarily propagates forward in time until becoming the state immediately before the final measurement (a quantum superposition of the problem-settings selected by the initial measurement each tensor product the corresponding solution). Time-symmetrically, the state selected by the final measurement (a single problem-setting tensor product the corresponding solution) propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous outcome of the initial measurement.
Of course, propagating backward in time the outcome of a measurement is unconventional. However, first, note that the outcome of the final measurement – a single problem-setting tensor product the corresponding solution – coincides with that of the ordinary quantum description of the quantum algorithm. As a consequence, mathematically (i.e. disregarding the time-direction of causality), its backward in time propagation, by the inverse of the time-forward unitary transformation, is the unitary part of the ordinary quantum description of the quantum algorithm back again. This backward in time propagation just undoes the sharing of the selection between initial and final measurements; it is therefore legitimate. Second, as we shall see, it is mandatory to complete the quantum description of the quantum algorithm in a way that describes the causal loops that it must host.
@@
We call the above sequence of operations the time-symmetrization of the ordinary quantum description of the quantum algorithm, because it is indeed the time-symmetrization of the time-reversible quantum process between the initial and final one-to-one correlated measurement outcomes.
@@
(i) each time symmetrization instance consists of a time-forward propagation (by a unitary transformation) going from the outcome of the initial measurement to the outcome of the final measurement and a time-backward propagation (by the inverse of the previous unitary transformation) going from the outcome of the latter measurement until replacing the previous outcome of the initial measurement. One can see that this is a causal loop that changes its own past.
(ii) the backward in time propagation of each time-symmetrization instance, which inherits both selections, is an instance of the time-symmetrized quantum algorithm. Since all backward in time propagations are the same, any one of them is the time-symmetrized quantum algorithm.
Summing up, we can see how the time-symmetrization process we are dealing with completes the ordinary (causal) quantum description of the quantum algorithm. First, it leaves the unitary part of the ordinary quantum description mathematically unaltered. Second, it changes the description of the behavior of causality along that part: the single causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
@@
Let us explain the reason for the specification “as viewed by the problem-solver”. The ordinary quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process. It cannot be with respect to the problem-solver: it would tell her, before she begins her problem-solving action, the problem-setting selected by the initial measurement and thus the solution of the problem (here, they are both the number of the drawer with the ball). To have the quantum description with respect to the problem-solver (relativized to her), we should conceal from her the problem-setting selected by the initial measurement (see subsection 2.1.2.)
Similarly, we should relativize to the problem-solver the causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, by hiding from her the part of the problem-setting selected by the initial measurement (see subsection 2.1.4).
Also note that the time-symmetrization we are dealing with is the only possible way of completing the ordinary quantum description of the quantum algorithm in a way that describes the causal loops logically implied by the sheer existence of its quantum speedup. In fact, to describe a causal loop (any one) we must properly share the selection of the problem-setting between initial and final measurements, with the part of the selection performed by the initial measurement propagating forward in time and that performed by the final measurement backward in time. To obtain exactly the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, this sharing must be even. This exactly identifies the time-symmetrization we have performed. It also shows that this time-symmetrization is mandatory: it is the only way of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup and that the quantum algorithm must host.
Further below we will show that all the above holds for a large class of quantum algorithms, those solving oracle problems. Note that, so completed, the quantum description of quantum algorithms says much more than their ordinary quantum description; indeed, it provides the following unified quantitative description of the quantum speedups:
@@
As is well known, in mainstream quantum computing science (which naturally relies on the notion of causal evolution), there is no unified quantitative description of the different quantum speedups. The only way of knowing the number of computation steps needed to solve a given oracle problem in an optimal quantum way is finding an optimal quantum algorithm that does that, naturally an object of great ingenuity and uncertain achievement.
Let us call the above statement in italics the advanced knowledge rule. This rule had already been achieved in the retrocausal approach to the quantum speedup of ourselves and coauthors $\left[4-8\right]$. In $\left[7\right]$, we had shown that it quantitatively explains the quantum speedup of all the major quantum algorithms. However, in these works, the time-symmetrization of the ordinary quantum description of the quantum algorithm was justified only by the fact that it turns out to produce, indeed, a unified quantitative description of the quantum speedups.
Now, this time-symmetrization has been shown to be necessary. It is the necessary consequence of the fact – we think inescapable – that the sheer existence of a quantum speedup over what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops that must be implicit in the quantum algorithm. This should be a significant advancement. Causal loops, hitherto foreign to quantum mechanics, would instead be common in it, at least as common as quantum algorithms are.
In Section 2.2 we will see that the above advanced knowledge rule generalizes in the following rule:
@@
In the first place, the rule in question says that the building up of quantum correlation between two initially unrelated observables can always occur with quantum speedup. Of course, the notion of advanced knowledge (on the part of the problem-solver) of half of the information about the solution would lose any meaning outside problem-solving. However, note that it can be replaced by that of a corresponding reduction of the dimension of the Hilbert space in which the search for one-to-one correlation between the two observables takes place. This search can therefore occur with quantum speedup (see Section 2.2 for further detail).
In the second place, the causal loop rule of quantum correlation also applies to the case that the unitary evolution between the two one-to-one correlated measurement outcomes does not build up any quantum correlation between the two observables because they are maximally entangled to start with. This is the case of quantum nonlocality, where there is one-to-one quantum correlation between the outcomes of measuring (in the same basis) two maximally entangled observables in a state of spatial separation between the two respective subsystems.
Here, the unitary evolution between the two one-to-one correlated measurement outcomes must go, first, backward in time from the outcome of the first measurement to the time when the two subsystems were not yet spatially separate and then (in causal order) forward in time to the outcome of the second measurement. Of course, the eigenvalue selected by either measurement is not changed by the unitary process of spatial separation between the two subsystems.
As shown in $\left[8\right]$, the causal loops generated by time-symmetrizing the ordinary quantum description of the unitary evolution in question also locally go from the initial to the final measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell the not yet measured observable the state to assume in view of its future measurement. Therefore, these causal loops exactly play the role of the hidden variables that Einstein and the others $\left[3\right]$ envisaged to locally explain quantum nonlocality. Note that, in Einstein’s deterministic view of the selection of a measurement outcome, whether the two one-to-one correlated measurement outcomes are determined by the initial or the final measurement, or by the two in an even way, is indifferent. In other words, the hidden variables explanation of quantum nonlocality suggested Einstein and the others is compatible with the present notion of the causal loops, the latter being the hidden variables those authors were looking for.
The above also says that the incompleteness that we are presently ascribing to the ordinary quantum description of quantum algorithms is exactly the same that Einstein and the others ascribed to that of quantum nonlocality. Of course, in both cases we are dealing with the incompleteness of the quantum description of a quantum correlation.
@@
As we have seen, completing the ordinary quantum description of quantum algorithms also highlights their teleological character, that is, their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again). We can say that, in the quantum world, we have found the missing physical basis for the notion of teleological evolution.
@@
This, naturally, raises the question whether the physical basis in question also applies to the evolutions of the living to which the teleological notion was once applied. The second part of the work addresses this question.
@@
Incidentally, being thought without physical basis, and specially because of its finalistic character, the teleological view is usually classified among the beliefs. This, however, should not conceal the fact that, to start with, it was motivated by obvious empirical reasons. Indeed, it fits the functional structures and the evolutions of the living much better than the causal view. Of course, besides being why it was conceived in the first place, this is why, in relatively recent times, whether to abandon it has long been debated among biologists (see subsection 3.4.).
@@
The question of whether the physical basis of the teleological character of quantum algorithms also applies to natural evolutions immediately raises a seemingly insurmountable problem. Quantum algorithms belong to the microscopic world; their implementations deal with a rather limited number of quantum bits and durations of a few tens microseconds. As things are now, larger sizes or durations soon bring to quantum decoherence and the process becomes classical. The evolutions of the living, instead, can be as large as the Earth and have durations of billions of years. They are therefore among the furthest things from the quantum processes of the microscopic world to which quantum algorithms belong.
@@
Of course, we are not discussing whether the chlorophyll cycle or orientation in migratory birds is quantum, which, as is well known, is considered to be a possibility. Here, the very enormity of the distance between the world of quantum algorithms and that of natural evolutions forbids any attempt to scale up toward the macroscopic world the quantum features of the microscopic one.